# CoTKR: Chain-of-Thought Enhanced Knowledge Rewriting for Complex Knowledge Graph Question Answering

### v3

@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable performance across various natural language processing tasks, marking a significant milestone Sanh et al. (2022); Brown et al. (2020); Zhang et al. (2022); Azaria et al. (2024); Chen (2024). Despite their superior performance in zero-shot scenarios Wei et al. (2022a); Kojima et al. (2022), they still encounter factual errors, known as “hallucinations” Ji et al. (2023b), especially in knowledge-intensive tasks like question answering (QA) Hu et al. (2023); Tan et al. (2023); Li et al. (2023); He et al. (2023). This issue arises due to the intrinsic limitations of LLMs, including factual inaccuracies and outdated knowledge Pan et al. (2023). To address this challenge, a substantial of work Ma et al. (2023a); Trivedi et al. (2023); Wu et al. (2023b) retrieves task-relevant knowledge from external sources as context, thereby enhancing the capabilities of LLMs in downstream tasks, known as Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Gao et al. (2023); Huang et al. (2023).
@@
Recent work Wu et al. (2023b); Baek et al. (2023); Sen et al. (2023); Wang et al. (2024) under the RAG paradigm explores the use of Knowledge Graphs (KGs) Pan et al. (2017b, a) as an information source to enhance the capabilities of LLMs in Question Answering (QA). Unlike typical QA tasks, a key challenge in KGQA under this paradigm lies in transforming question-related subgraphs into natural language that LLMs can understand while preserving the structural information Ko et al. (2024); Ding et al. (2024); Wu et al. (2023b). This process is referred to as Knowledge Rewriting (KR) in this study. As illustrated in Figure 1, this paper summarizes the commonly used knowledge rewriting methods in existing work. Most prior studies Baek et al. (2023); Sen et al. (2023); Wang et al. (2023a) employ simple linear concatenation method (Triple), which concatenates the subject, relation, and object of each triple to form triple-form text. Additionally, considering that LLMs are pre-trained on text corpora and struggle with structured triple-form text, some efforts Wu et al. (2023b); Bian et al. (2021); Chen et al. (2022) focus on converting triples into natural language through KG-to-Text. Furthermore, given that retrieved subgraphs often contain redundant information irrelevant to the question, other studies Ko et al. (2024); Dernbach et al. (2024) aim to extract question-relevant knowledge from the triples to generate summary pertinent to the question.
@@
To this end, we propose Chain-of-Thought Enhanced Knowledge Rewriting, CoTKR. Inspired by ReAct Yao et al. (2023), the core of our method involves generating reasoning traces and corresponding knowledge in an interleaved manner. As shown in Figure 1, we alternate the following two operations: (1) Reasoning: decomposing the question to identify the knowledge required for inference; (2) Summarization: summarizing the relevant knowledge from retrieved triples, informed by the reasoning step’s output. By integrating Chain-of-Thought (CoT) Wei et al. (2022b) with knowledge rewriting, CoTKR filters out irrelevant information and extracts question-related knowledge. Moreover, it generates a well-organized knowledge representation [^2] In this paper, “knowledge representation” refers to the natural language form of question-related knowledge. semantically aligned with the question. Unlike traditional CoT applications in QA, our framework employs the knowledge rewriter to first summarize knowledge, which then serves as contextual information to enhance QA performance. This strategy offers superior robustness. Although the summary might be inaccurate, it still contributes valuable information, potentially leading to correct answers. However, applying CoT to QA requires more precise reasoning chains, which are significantly affected by the error propagation Wang et al. (2023a); Yao et al. (2023). To train knowledge rewriters based on LLMs, we design a training framework for CoTKR. In the first stage, inspired by previous work Ma et al. (2023a); Wu et al. (2023b); Ko et al. (2024), we use knowledge representations generated by ChatGPT to guide the supervised fine-tuning of the knowledge rewriter, enabling it to initially master the capability of knowledge rewriting. In the second stage, we introduce Preference Alignment from Question Answering Feedback (PAQAF) to bridge the preference gap between the knowledge rewriter and the QA model. This method evaluates the quality of different knowledge representations based on the corresponding responses from the QA model. Subsequently, it constructs preference pairs, and fine-tunes LLMs through direct preference optimization (DPO) Rafailov et al. (2023).
We conduct experiments on GrailQA Gu et al. (2021) and GraphQuestions Su et al. (2016), comparing commonly used knowledge rewriting methods in existing work. Contrary to previous findings Dai et al. (2024); Baek et al. (2023), which suggest that LLMs perform better with knowledge in triple-form rather than in natural language, our findings demonstrate that LLMs can significantly benefit from knowledge represented in carefully crafted natural language. This indicates that our method could substantially enhance the performance of LLMs in KGQA.
@@ 2.1 KG-Augmented LLMs for KGQA
To mitigate hallucination in LLMs, existing work Wu et al. (2023b); Baek et al. (2023); Sen et al. (2023); Wang et al. (2024) attempts to enhance LLMs with KGs in the RAG paradigm. The naïve approach involves retrieving question-related triples from KGs as contextual information for QA Baek et al. (2023); Sen et al. (2023). Although this method has proven effective, there remains ample room for improvement. Some studies Wang et al. (2024, 2023a) integrate Chains-of-Thought (CoT) Wei et al. (2022b) with RAG to tackle complex questions. KeqingWang et al. (2024) decomposes complex questions using predefined templates, retrieves candidate entities from KG, reasons through sub-questions, and ultimately generates answers with clear reasoning paths. KD-CoT Wang et al. (2023a) validates and adjusts reasoning traces in CoT through interactions with external knowledge, thereby addressing issues of hallucinations and error propagation. Furthermore, alternative efforts Wu et al. (2023b); Ko et al. (2024) address the limitations of LLMs in processing structured knowledge and the noise in retrieved triples by post-processing these triples into natural language or summaries pertinent to the questions.
@@ 2.2 Preference Alignment for LLMs on Question Answering
LLMs have the potential to generate content that contains gender discrimination, unethical elements, and racial biases, inconsistent with human values Wu et al. (2023a); Ray (2023). To address this issue, Preference Alignment (PA) Ji et al. (2023a); Wang et al. (2023b) aims to fine-tune LLMs to align with human preferences. Existing QA work based on LLMs uses PA to bridge the gap between model preferences and those of humans or the QA tasks. KnowPAT Zhang et al. (2023) trains LLMs on a knowledge preference set to align their knowledge biases with human preferences, selecting better factual knowledge as context. BGM Ke et al. (2024) utilizes downstream task metrics as rewards to optimize the bridging model between retrievers and QA models. Rewrite-Retrieve-Read Ma et al. (2023b) employs QA evaluation metrics as reward signals, fine-tuning the query rewriting module. EFSUM Ko et al. (2024) constructs preference pairs sampled from LLMs and fine-tunes the knowledge summarizer using the Direct Preference Optimization (DPO) Rafailov et al. (2023) algorithm.
@@ 3 Preliminaries
Knowledge Graph Question Answering (KGQA) aims to answer natural language questions by utilizing a set of facts within KGs. Following previous work Saxena et al. (2020); Jiang et al. (2023b), we assume that the subject entity of the question is given. Given a question $q$ and a subject entity $e$, the objective is to generate a response $a$ using the relevant facts in the KG $G$ that accurately addresses the question.
@@ 4.1 Chain-of-Thought Enhanced Knowledge Rewriting
Assume we have the reasoning traces at step $t-1$ as $x_{t-1,r}$ and the summarized knowledge at step $t-1$ as $x_{t-1,k}$. The corresponding knowledge representation, i.e., $X_{t-1}$ is represented as:
@@ 4.2.1 Supervised Fine-tuning with Knowledge Distilled from ChatGPT
Reference Knowledge Representation Generation. Inspired by previous work Ma et al. (2023a); Wu et al. (2023b); Ko et al. (2024), we employ ChatGPT as the data generator to construct training corpora. We verbalize the question-related subgraph, $G^{\prime}$, through simple linear concatenation and combine it with the question, $q$, to form the input prompt $x$. Subsequently, ChatGPT generates the reference knowledge representation $k$ based on several examples (i.e., demonstrations) and the provided input $x$. Finally, we construct the training dataset $D_{T}=\{(x_{1},k_{1}),(x_{2},k_{2}),...,(x_{T},k_{T})\}$.
@@ 4.2.2 Preference Alignment from Question Answering Feedback (PAQAF)
Preference Annotation based on Question Answering Feedback. Among the candidate knowledge representations, we select the two, $k_{1}$ and $k_{2}$, with the greatest semantic difference (i.e., the lowest similarity) to facilitate faster convergence during training. Utilizing standard evaluation methods for assessing these representations is suboptimal, as they fail to align with the preferences of QA models. Inspired by the findings in previous workWu et al. (2023b); Ko et al. (2024); Zhang et al. (2023), we posit that better knowledge representations generally lead to better performance on QA. Consequently, we adopt $k_{1}$ and $k_{2}$ as contextual knowledge, prompting the QA model $Q$ to answer the question $q$, generating answers $a_{1}$ and $a_{2}$, respectively. Subsequently, we prompt ChatGPT to assess the quality of $a_{1}$ and $a_{2}$ from the perspectives of accuracy and relevance. This evaluation aims to identify the preferred knowledge representation $k^{+}$ and the dispreferred knowledge representation $k^{-}$. Details of the evaluation prompt are provided in Appendix 8.
@@ 5.1 Datasets
GrailQA Gu et al. (2021) is a challenging, large-scale multi-hop KGQA benchmark that consists of 64,331 questions (44,337 train, 6,763 dev, 13,231 test). The training and dev sets provide annotated SPARQL query and answer entities, while the test set comprises only the questions. For evaluation convenience, the dev set is used for testing.
GraphQuestions Su et al. (2016) is a characteristic-rich dataset for factoid question answering based on Freebase. It comprises 5,166 questions (2,771 train, 2,395 test). For each question, the dataset provides corresponding SPARQL query and answer entities.
@@ 5.2 Large Language Models
In this experiment, Llama-2 (7B) Touvron et al. (2023b), Llama-3 (8B) AI@Meta (2024), and ChatGPT  [^3] https://api.openai.com/ are employed for knowledge rewriting, while ChatGPT and Mistral (7B) Jiang et al. (2023a) are used for QA tasks. The details of these LLMs are provided in Appendix A.3.
@@ 5.3 Baselines
Simple linear concatenation (Triple) Baek et al. (2023); Sen et al. (2023) concatenates the subject, predicate, and object of a triple to generate triple-form text. This method does not require additional models for knowledge rewriting.
KG-to-Text Wu et al. (2023b) transforms facts into the free-form text for each relation path with a KG-to-Text model, addressing the limitations of LLMs in understanding structured triple-form text.
Summary Ko et al. (2024) converts triples into a question-relevant summary, alleviating the issue of redundant contextual knowledge.
@@ 5.4 Retrieval Methods
BM25. We follow the processing method in DecAF Yu et al. (2023), simply linearizing the 1-hop subgraph of the topic entity as the article. We take the top 30 triples corresponding to the candidate documents as the retrieval results.
@@ 5.5 Evaluation Metrics
Following previous work on generative KGQA Wu et al. (2023b); Baek et al. (2023); Ko et al. (2024), we adopt Accuracy (Acc) as one of our evaluation metrics. It measures whether the model’s response includes at least one answer entity. For a dataset comprising $N$ questions, Acc is calculated as follows:
@@
$Acc_{i}=\begin{cases}1,&\text{if at least one answer entity}\\ &\text{appears in the response}\\ 0,&\text{if no answer entity appears in}\\ &\text{the response}\end{cases}$ (8)
@@ 5.9 Effectiveness of Data Augmentation
[…]
@@
[…]
@@ Acknowledgments
This work is partially supported by National Nature Science Foundation of China under No. U21A20488, by the project "Key Laboratory of rich-media Digital Publishing Content Organization and Knowledge Service Open Fund-Research on Knowledge-enhanced Training Techinques of Large Language Model" No. ZD2024-04/01, by the EPSRC project OntoEm (EP/Y017706/1), and by Southeast University-China Mobile Research Institute Joint Innovation Center. We thank the Big Data Computing Center of Southeast University for providing the facility support on the numerical calculations in this paper. ChatGPT was used to enhance the readability of some of the text and improve the language of this paper, after the content was first added manually. All material was then checked manually before submission.

### v4

@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable performance across various natural language processing tasks, marking a significant milestone DBLP:conf/iclr/SanhWRBSACSRDBX22; DBLP:conf/nips/BrownMRSKDNSSAA20; DBLP:journals/corr/abs-2205-01068; DBLP:journals/dint/AzariaAR24; publisher/Beijing. Despite their superior performance in zero-shot scenarios DBLP:conf/iclr/WeiBZGYLDDL22; DBLP:conf/nips/KojimaGRMI22, they still encounter factual errors, known as “hallucinations” DBLP:journals/csur/JiLFYSXIBMF23, especially in knowledge-intensive tasks like question answering (QA) DBLP:journals/www/HuWQMCPA23; DBLP:journals/corr/abs-2303-07992; DBLP:journals/dint/LiZLYC23; HUGP2023. This issue arises due to the intrinsic limitations of LLMs, including factual inaccuracies and outdated knowledge PRKSC2023. To address this challenge, a substantial of work DBLP:conf/emnlp/MaGHZD23; DBLP:conf/acl/TrivediBKS23; DBLP:journals/corr/abs-2309-11206 retrieves task-relevant knowledge from external sources as context, thereby enhancing the capabilities of LLMs in downstream tasks, known as Retrieval-Augmented Generation (RAG) DBLP:conf/nips/LewisPPPKGKLYR020; DBLP:journals/corr/abs-2312-10997; HLVPP2023.
@@
Recent work DBLP:journals/corr/abs-2309-11206; baek-etal-2023-knowledge; sen-etal-2023-knowledge; DBLP:journals/corr/abs-2401-00426 under the RAG paradigm explores the use of Knowledge Graphs (KGs) PVGW2017; PCEH+2017 as an information source to enhance the capabilities of LLMs in Question Answering (QA). Unlike typical QA tasks, a key challenge in KGQA under this paradigm lies in transforming question-related subgraphs into natural language that LLMs can understand while preserving the structural information DBLP:journals/corr/abs-2403-02966; DBLP:conf/www/DingLLQ24; DBLP:journals/corr/abs-2309-11206. This process is referred to as Knowledge Rewriting (KR) in this study. As illustrated in Figure 1, this paper summarizes the commonly used knowledge rewriting methods in existing work. Most prior studies baek-etal-2023-knowledge; sen-etal-2023-knowledge; DBLP:journals/corr/abs-2308-13259 employ simple linear concatenation method (Triple), which concatenates the subject, relation, and object of each triple to form triple-form text. Additionally, considering that LLMs are pre-trained on text corpora and struggle with structured triple-form text, some efforts DBLP:journals/corr/abs-2309-11206; DBLP:conf/aaai/BianH0021; DBLP:conf/jist/0007HCGFP0Z22 focus on converting triples into natural language through KG-to-Text. Furthermore, given that retrieved subgraphs often contain redundant information irrelevant to the question, other studies DBLP:journals/corr/abs-2403-02966; DBLP:conf/aaaiss/DernbachAZHC24 aim to extract question-relevant knowledge from the triples to generate summary pertinent to the question.
@@
To this end, we propose Chain-of-Thought Enhanced Knowledge Rewriting, CoTKR. Inspired by ReAct DBLP:conf/iclr/YaoZYDSN023, the core of our method involves generating reasoning traces and corresponding knowledge in an interleaved manner. As shown in Figure 1, we alternate the following two operations: (1) Reasoning: decomposing the question to identify the knowledge required for inference; (2) Summarization: summarizing the relevant knowledge from retrieved triples, informed by the reasoning step’s output. By integrating Chain-of-Thought (CoT) DBLP:conf/nips/Wei0SBIXCLZ22 with knowledge rewriting, CoTKR filters out irrelevant information and extracts question-related knowledge. Moreover, it generates a well-organized knowledge representation [^2] In this paper, “knowledge representation” refers to the natural language form of question-related knowledge. semantically aligned with the question. Unlike traditional CoT applications in QA, our framework employs the knowledge rewriter to first summarize knowledge, which then serves as contextual information to enhance QA performance. This strategy offers superior robustness. Although the summary might be inaccurate, it still contributes valuable information, potentially leading to correct answers. However, applying CoT to QA requires more precise reasoning chains, which are significantly affected by the error propagation DBLP:journals/corr/abs-2308-13259; DBLP:conf/iclr/YaoZYDSN023. To train knowledge rewriters based on LLMs, we design a training framework for CoTKR. In the first stage, inspired by previous work DBLP:conf/emnlp/MaGHZD23; DBLP:journals/corr/abs-2309-11206; DBLP:journals/corr/abs-2403-02966, we use knowledge representations generated by ChatGPT to guide the supervised fine-tuning of the knowledge rewriter, enabling it to initially master the capability of knowledge rewriting. In the second stage, we introduce Preference Alignment from Question Answering Feedback (PAQAF) to bridge the preference gap between the knowledge rewriter and the QA model. This method evaluates the quality of different knowledge representations based on the corresponding responses from the QA model. Subsequently, it constructs preference pairs, and fine-tunes LLMs through direct preference optimization (DPO) DBLP:conf/nips/RafailovSMMEF23.
We conduct experiments on GrailQA DBLP:conf/www/GuKVSLY021 and GraphQuestions DBLP:conf/emnlp/SuSSSGYY16, comparing commonly used knowledge rewriting methods in existing work. Contrary to previous findings DBLP:journals/corr/abs-2402-11541; baek-etal-2023-knowledge, which suggest that LLMs perform better with knowledge in triple-form rather than in natural language, our findings demonstrate that LLMs can significantly benefit from knowledge represented in carefully crafted natural language. This indicates that our method could substantially enhance the performance of LLMs in KGQA.
@@ 2.1 KG-Augmented LLMs for KGQA
To mitigate hallucination in LLMs, existing work DBLP:journals/corr/abs-2309-11206; baek-etal-2023-knowledge; sen-etal-2023-knowledge; DBLP:journals/corr/abs-2401-00426 attempts to enhance LLMs with KGs in the RAG paradigm. The naïve approach involves retrieving question-related triples from KGs as contextual information for QA baek-etal-2023-knowledge; sen-etal-2023-knowledge. Although this method has proven effective, there remains ample room for improvement. Some studies DBLP:journals/corr/abs-2401-00426; DBLP:journals/corr/abs-2308-13259 integrate Chains-of-Thought (CoT) DBLP:conf/nips/Wei0SBIXCLZ22 with RAG to tackle complex questions. KeqingDBLP:journals/corr/abs-2401-00426 decomposes complex questions using predefined templates, retrieves candidate entities from KG, reasons through sub-questions, and ultimately generates answers with clear reasoning paths. KD-CoT DBLP:journals/corr/abs-2308-13259 validates and adjusts reasoning traces in CoT through interactions with external knowledge, thereby addressing issues of hallucinations and error propagation. Furthermore, alternative efforts DBLP:journals/corr/abs-2309-11206; DBLP:journals/corr/abs-2403-02966 address the limitations of LLMs in processing structured knowledge and the noise in retrieved triples by post-processing these triples into natural language or summaries pertinent to the questions.
@@ 2.2 Preference Alignment for LLMs on Question Answering
LLMs have the potential to generate content that contains gender discrimination, unethical elements, and racial biases, inconsistent with human values DBLP:journals/corr/abs-2307-14192; ray2023chatgpt. To address this issue, Preference Alignment (PA) DBLP:journals/corr/abs-2310-19852; DBLP:journals/corr/abs-2307-12966 aims to fine-tune LLMs to align with human preferences. Existing QA work based on LLMs uses PA to bridge the gap between model preferences and those of humans or the QA tasks. KnowPAT DBLP:journals/corr/abs-2311-06503 trains LLMs on a knowledge preference set to align their knowledge biases with human preferences, selecting better factual knowledge as context. BGM DBLP:journals/corr/abs-2401-06954 utilizes downstream task metrics as rewards to optimize the bridging model between retrievers and QA models. Rewrite-Retrieve-Read ma-etal-2023-query employs QA evaluation metrics as reward signals, fine-tuning the query rewriting module. EFSUM DBLP:journals/corr/abs-2403-02966 constructs preference pairs sampled from LLMs and fine-tunes the knowledge summarizer using the Direct Preference Optimization (DPO) DBLP:conf/nips/RafailovSMMEF23 algorithm.
@@ 3 Preliminaries
Knowledge Graph Question Answering (KGQA) aims to answer natural language questions by utilizing a set of facts within KGs. Following previous work DBLP:conf/acl/SaxenaTT20; DBLP:conf/iclr/JiangZ0W23, we assume that the subject entity of the question is given. Given a question $q$ and a subject entity $e$, the objective is to generate a response $a$ using the relevant facts in the KG $G$ that accurately addresses the question.
@@ 4.1 Chain-of-Thought Enhanced Knowledge Rewriting
Assume we have the reasoning traces at step t1t-1 as $x_{t-1,r}$ and the summarized knowledge at step t1t-1 as $x_{t-1,k}$. The corresponding knowledge representation, i.e., $X_{t-1}$ is represented as:
@@ 4.2.1 Supervised Fine-tuning with Knowledge Distilled from ChatGPT
Reference Knowledge Representation Generation. Inspired by previous work DBLP:conf/emnlp/MaGHZD23; DBLP:journals/corr/abs-2309-11206; DBLP:journals/corr/abs-2403-02966, we employ ChatGPT as the data generator to construct training corpora. We verbalize the question-related subgraph, $G^{\prime}$, through simple linear concatenation and combine it with the question, $q$, to form the input prompt $x$. Subsequently, ChatGPT generates the reference knowledge representation $k$ based on several examples (i.e., demonstrations) and the provided input $x$. Finally, we construct the training dataset $D_{T}=\{(x_{1},k_{1}),(x_{2},k_{2}),...,(x_{T},k_{T})\}$.
@@ 4.2.2 Preference Alignment from Question Answering Feedback (PAQAF)
Preference Annotation based on Question Answering Feedback. Among the candidate knowledge representations, we select the two, $k_{1}$ and $k_{2}$, with the greatest semantic difference (i.e., the lowest similarity) to facilitate faster convergence during training. Utilizing standard evaluation methods for assessing these representations is suboptimal, as they fail to align with the preferences of QA models. Inspired by the findings in previous workDBLP:journals/corr/abs-2309-11206; DBLP:journals/corr/abs-2403-02966; DBLP:journals/corr/abs-2311-06503, we posit that better knowledge representations generally lead to better performance on QA. Consequently, we adopt $k_{1}$ and $k_{2}$ as contextual knowledge, prompting the QA model $Q$ to answer the question $q$, generating answers $a_{1}$ and $a_{2}$, respectively. Subsequently, we prompt ChatGPT to assess the quality of $a_{1}$ and $a_{2}$ from the perspectives of accuracy and relevance. This evaluation aims to identify the preferred knowledge representation $k^{+}$ and the dispreferred knowledge representation $k^{-}$. Details of the evaluation prompt are provided in Appendix 8.
@@ 5.1 Datasets
GrailQA DBLP:conf/www/GuKVSLY021 is a challenging, large-scale multi-hop KGQA benchmark that consists of 64,331 questions (44,337 train, 6,763 dev, 13,231 test). The training and dev sets provide annotated SPARQL query and answer entities, while the test set comprises only the questions. For evaluation convenience, the dev set is used for testing.
GraphQuestions DBLP:conf/emnlp/SuSSSGYY16 is a characteristic-rich dataset for factoid question answering based on Freebase. It comprises 5,166 questions (2,771 train, 2,395 test). For each question, the dataset provides corresponding SPARQL query and answer entities.
@@ 5.2 Large Language Models
In this experiment, Llama-2 (7B) touvron2023llama2, Llama-3 (8B) llama3modelcard, and ChatGPT  [^3] https://api.openai.com/ are employed for knowledge rewriting, while ChatGPT and Mistral (7B) jiang2023mistral are used for QA tasks. The details of these LLMs are provided in Appendix A.3.
@@ 5.3 Baselines
Simple linear concatenation (Triple) baek-etal-2023-knowledge; sen-etal-2023-knowledge concatenates the subject, predicate, and object of a triple to generate triple-form text. This method does not require additional models for knowledge rewriting.
KG-to-Text DBLP:journals/corr/abs-2309-11206 transforms facts into the free-form text for each relation path with a KG-to-Text model, addressing the limitations of LLMs in understanding structured triple-form text.
Summary DBLP:journals/corr/abs-2403-02966 converts triples into a question-relevant summary, alleviating the issue of redundant contextual knowledge.
@@ 5.4 Retrieval Methods
BM25. We follow the processing method in DecAF DBLP:conf/iclr/YuZNZL0HWWX23, simply linearizing the 1-hop subgraph of the topic entity as the article. We take the top 30 triples corresponding to the candidate documents as the retrieval results.
@@ 5.5 Evaluation Metrics
Following previous work on generative KGQA DBLP:journals/corr/abs-2309-11206; baek-etal-2023-knowledge; DBLP:journals/corr/abs-2403-02966, we adopt Accuracy (Acc) as one of our evaluation metrics. It measures whether the model’s response includes at least one answer entity. For a dataset comprising $N$ questions, Acc is calculated as follows:
@@
[…]
@@ 5.9 Effectiveness of Data Augmentation
CoTKR Reason: I need to know the unit of area associated with the measurement system that has an electric field strength unit of Volt per meter.
@@
CoTKR+PA Reason: I need to know which measurement system includes the unit for electric field strength in Volt per meter.
@@ Acknowledgments
This work is partially supported by National Nature Science Foundation of China under No. U21A20488, by the project "Key Laboratory of rich-media Digital Publishing Content Organization and Knowledge Service Open Fund-Research on Knowledge-enhanced Training Techinques of Large Language Model" No. ZD2024-04/01, by the EPSRC project OntoEm (EP/Y017706/1), and by Southeast University-China Mobile Research Institute Joint Innovation Center. It is funded by Southeast University-China Mobile Research Institute Joint Innovation Center. We thank the Big Data Computing Center of Southeast University for providing the facility support on the numerical calculations in this paper. ChatGPT was used to enhance the readability of some of the text and improve the language of this paper, after the content was first added manually. All material was then checked manually before submission.